The physician orders gr 10 of Tylenol suppository PR Q6H for a patient who will be transferred back to an assisted living facility. The clinic has 650mg suppositories on hand. How many suppositories would you administer per dose?

You administer 10 mL of Mylanta to a patient before they leave to return home. Since the patient will receive the rest of the doses at home you will need to explain the dose in household measurements, how many teaspoons of Mylanta will the patient need to take?
Answer #1

ANSWER.

* Calculation of number of suppositories required per dose.

           Ordered dose = gr 10.

      Convert gr to mg for calculation

               1 gr = 64.80 mg.

     So

           Ordered dose = 64.80 * 10

                                   = 648 mg

         Available dose = 650 mg/suppository.

Number of suppositories required = Ordered dose/Available dose.

                                                         = 648/650

                                                         = 0.99

                                                         = 1 suppository per dose.

* Calculation of teaspoon of drug required.

           Volume(ml) of drug required for administration = 10 ml.

     Convert ml to teaspoon for calculation

                       1 teaspoon = 5 ml.

   So

       Teaspoon of drug required = 10/5

                                                   = 2 teaspoon.